OpenStudy (anonymous):

Solve the quantity 2 x minus 5 divided by the quantity x minus 2, minus 2 equals 3 divided by the quantity x plus 2

OpenStudy (anonymous):

2x-5 3 ----- -2 = ---- x-2 x+2 2x-5-2x+4 3 ---------- = --- x-2 x+2 -1 3 ----- = ------ x-2 x+2 now do cross multiplication -1(x+2) = 3(x-2) -x-2 = 3x-6 3x+x =6 +2 4x=8 x=8/4 x=2

OpenStudy (anonymous):

@Farheen28 check 4x=4 in your sol.

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

what you want to ask?

OpenStudy (anonymous):

check your steps

OpenStudy (anonymous):

x will be 1

OpenStudy (anonymous):

oops yes you are right :)

OpenStudy (anonymous):

2x-5 3 ----- -2 = ---- x-2 x+2 2x-5-2x+4 3 ---------- = --- x-2 x+2 -1 3 ----- = ------ x-2 x+2 now do cross multiplication -1(x+2) = 3(x-2) -x-2 = 3x-6 3x+x =6 -2 4x=4 x=1

OpenStudy (anonymous):

when you do cross multiplication it means you are eliminating x = 2 frm solution

OpenStudy (anonymous):

thanks :) i don't know where my mind was :P